apter 10 Quiz 6 Seved Help Save & 4 Piper Corporations standards call for 3,750 direct labor-hours to produce 1,250 units of product. During October the company worked 950 direct labor-hours and produced 950 units. The standard hours allowed for October would be: 00:57:55 Multiple Choice 3,750 hours 150 hours 2,850 hours 2,800 hours
Answer #1

Standards hour per Unit of Output = 3750/1250 = 3

Standards Hours Allowed = 3*950 =2850 hours

Option C is Correct